# Real-World Adversarial Attacks on RF-Based Drone Detectors

### v2

@@ Abstract
Index Terms—  Drone detection, Adversarial ML., Physical Attack, RF Sensing, Over-the-Air Attack, I/Q domain
@@ 4.1 Dataset
Attack set generation. We assume an adversary that aims to prevent their drone’s detection by OD systems. They can control their drone’s transmit channel (i.e., center frequency) but have no control of other drones’ transmission channels. Accordingly, we recorded the target drone OTA in a realistic, noisy environment; other drones’ signals were digitally mixed into the I/Q stream at random channels across the 2.4 GHz ISM band. This setup reflects a realistic threat model in which the adversary can manipulate the target drone’s emissions via a nearby adversarial transmitter but cannot influence coexisting signals. The attack set used to train the adversarial waveform was split into 70%, 15%, and 15% for the training, validation, and test subsets, respectively.
@@ 4.4 Experimental Results
Table 1: The AP (for SPR = 10 dB) in the following settings: clean, AWGN, WB, GB, CS, and $B$. $\downarrow$ indicates that lower values are better and $\uparrow$ indicates that higher values are better.
@@
Physical attacks. Due to space limitations, we present only the closed-set CUAP, which exhibited the strongest cross-architecture transfer in digital evaluations and represents a practical scenario when the target detector architecture is unknown, demonstrating OTA feasibility in a realistic multi-emitter environment. Drones were positioned at fixed distances (3m, 5m, and 7m) from the receiving antenna to assess range effects. Since OTA transmissions timing and phase are non-repeatable, Clean and Attacked settings use separate recordings acquired consecutively under identical setup.
The results presented in Table 2 highlight the transferability of our digital optimization approach to physical attacks. Under attack, the target class MDR reached near-complete suppression ($\geq 99\%$) on most models across distances, with YOLOv9 at 3m and FRCNN as exceptions. Detection of the non-target classes remained largely preserved (MDR degradation $\leq 18.3\%$), confirming a class-specific targeting effect. These OTA results demonstrate that I/Q domain perturbations retain adversarial effects after propagation through physical RF channels across varying ranges.
@@
Table 2: CUAP (CS) OTA performance at different distances. High MDR for the target class indicates a successful attack. $\downarrow$ indicates that lower values are better and $\uparrow$ indicates that higher values are better.

### v3

@@ Abstract
Index Terms—  Drone detection, Adversarial ML.
@@ 4.1 Dataset
Attack set generation. We assume an adversary that aims to prevent their drone’s detection by OD systems. They can control their drone’s transmit channel (i.e., center frequency) but have no control of other drones’ transmission channels. Accordingly, we recorded the target drone OTA in a realistic, noisy environment; other drones’ signals were digitally mixed into the I/Q stream at random channels across the 2.4 GHz ISM band. This setup reflects a realistic threat model in which the adversary can manipulate the target drone’s emissions but cannot influence coexisting signals. The attack set used to train the adversarial waveform was split into 70%, 15%, and 15% for the training, validation, and test subsets, respectively.
@@ 4.4 Experimental Results
Table 1: AP under digital attacks (for SPR = 10 dB) in the following settings: clean, AWGN, WB, GB, CS, and $B$.
@@
Physical attacks. Due to space limitations, we present only the closed-set CUAP, which exhibited the strongest cross-architecture transfer in digital evaluations to demonstrate OTA feasibility in a realistic multi-emitter environment. Following the described threat model, the target drone operated in the same predetermined channel, while the other three drones independently selected transmission channels, eliminating spectral hopping and isolating OTA effects. The optimized $\delta$ was transmitted alongside legitimate communications, with transmit power adjusted to match the received perturbation levels from digital experiments. Drones were placed at fixed distances (3 m, 5 m, and 7 m) from the receiving antenna to assess range effects. Clean and Attacked recordings were acquired consecutively under identical conditions due to non-repeatable OTA timing and phase.
The results presented in Table 2 highlight the transferability of our digital optimization approach to a physical attack designed for the Mavic 2 Zoom. The target class MDR reached near-complete suppression ($\geq 99\%$) on most models across distances, with YOLOv9 at 3 m and FRCNN as exceptions. Detection of the non-target classes remained largely preserved (MDR degradation $\leq 18.3\%$), confirming a class-specific targeting effect. These OTA results demonstrate that I/Q domain perturbations retain adversarial effects after propagation through physical RF channels across varying ranges.
@@
Table 2: OTA performance (CS setting, Mavic 2 Zoom) at varying distances. High MDR for the target class indicates successful attack.